# Predicting Chess Puzzle Difficulty with Transformers

### v1

@@ Abstract
Experimental evaluation was conducted on a dataset of over 4 million chess puzzles. Results demonstrate GlickFormer’s superior performance compared to the state-of-the-art ChessFormer baseline across multiple metrics. The algorithm’s performance has also been recognized through its competitive results in the IEEE BigData 2024 Cup: Predicting Chess Puzzle Difficulty competition. This paper has been submitted to 2024 IEEE International Conference on Big Data and is currently under review.
@@ II Related Work
In a recent development, a new study has emerged [13], which participated in the same IEEE BigData 2024 Cup competition as ours. It employs a combination of Convolutional Neural Networks (CNN) and bidirectional Long Short-Term Memory (LSTM) networks. While our approach leverages transformer architecture, this study opts for more traditional deep learning methods, while still incorporating temporal information. The CNN-LSTM approach achieved comparable, albeit lower, results on the public leaderboard in both preliminary and final scores, compared to our transformer-based solution. This parallel development in the field underscores the complexity of chess rating estimation and the diversity of potential approaches. While the CNN-LSTM model demonstrates the viability of traditional deep learning methods for this task, the superior performance of our transformer-based architecture suggests that the self-attention mechanism may be particularly well-suited for capturing the nuanced patterns in chess problem difficulty.
@@ V-A Experimental Setup
The models were trained for 11.000 steps using the RMSprop optimizer with a learning rate of $1\times 10^{-6}$, $\rho=0.99$ and a batch size of 4096. We observed that momentum-based optimizers like Adam or higher learning rates caused the models to converge prematurely to predicting the mean of the dataset with zero standard deviation.
@@ 3. Accuracy within 1, 2, and 3 Rating Deviations:
ChessFormer 277.20 3.28
Factorized Self-Attention GlickFormer 265.87 3.135
Factorized Encoder GlickFormer 258.72 3.062
@@
ChessFormer 20.16 38.48 54.58
Factorized Self-Attention GlickFormer 21.31 40.53 56.90
Factorized Encoder GlickFormer 21.57 41.31 57.81
@@
The Mean Absolute Error (MAE) and Mean Absolute Z-Score (MAZ) are lower for our models, indicating more accurate predictions. Specifically, the Factorized Encoder GlickFormer achieves the lowest MAE of 258.72, demonstrating the effectiveness of separating spatial and temporal processing.
The accuracy within 1, 2, and 3 rating deviations also shows improvement with our models. The Factorized Encoder GlickFormer achieves the highest accuracy within all deviation ranges, with 57.81% of predictions within 3 rating deviations of the true ratings.

### v2

@@ Abstract
Experimental evaluation was conducted on a dataset of over 4 million chess puzzles. Results demonstrate GlickFormer’s superior performance compared to the state-of-the-art ChessFormer baseline across multiple metrics. The algorithm’s performance has also been recognized through its competitive results in the IEEE BigData 2024 Cup: Predicting Chess Puzzle Difficulty competition, where it placed 11th.
@@ II Related Work
In a recent development, a new study has emerged [13], which participated in the same IEEE BigData 2024 Cup competition [14] as ours. It employs a combination of Convolutional Neural Networks (CNN) and bidirectional Long Short-Term Memory (LSTM) networks. While our approach leverages transformer architecture, this study opts for more traditional deep learning methods, while still incorporating temporal information. The CNN-LSTM approach achieved comparable, albeit lower, results on the public leaderboard in both preliminary and final scores, compared to our transformer-based solution. This parallel development in the field underscores the complexity of chess rating estimation and the diversity of potential approaches. While the CNN-LSTM model demonstrates the viability of traditional deep learning methods for this task, the superior performance of our transformer-based architecture suggests that the self-attention mechanism may be particularly well-suited for capturing the nuanced patterns in chess problem difficulty.
@@ V-A Experimental Setup
The models were trained for 28.000 steps using the RMSprop optimizer with a learning rate of $1\times 10^{-6}$, $\rho=0.99$ and a batch size of 4096. We observed that momentum-based optimizers like Adam or higher learning rates caused the models to converge prematurely to predicting the mean of the dataset with zero standard deviation. Additionally, we employed a cyclical optimizer state restarting strategy [19], where the length of the $k$-th cycle was set to $1000k$ steps.
@@ 3. Accuracy within 1, 2, and 3 Rating Deviations:
ChessFormer 227.00 2.68
Factorized Self-Attention GlickFormer 221.80 2.62
Factorized Encoder GlickFormer 217.71 2.57
@@
ChessFormer 25.42 48.21 65.29
Factorized Self-Attention GlickFormer 26.13 48.54 66.08
Factorized Encoder GlickFormer 27.00 50.45 67.66
@@
The Mean Absolute Error (MAE) and Mean Absolute Z-Score (MAZ) are lower for our models, indicating more accurate predictions. Specifically, the Factorized Encoder GlickFormer achieves the lowest MAE of 217.71, demonstrating the effectiveness of separating spatial and temporal processing.
The accuracy within 1, 2, and 3 rating deviations also shows improvement with our models. The Factorized Encoder GlickFormer achieves the highest accuracy within all deviation ranges, with 67.66% of predictions within 3 rating deviations of the true ratings.